JUUlDS of medication will you draw up? PEEL Use the following information for questions 23 -24. Order: Augmentin 5mg/kgPO q6hNot sure how to do #24.
Answer #1

Order of Augmentin is 5mg/kgPOq6h

patint weight 41 lbs

Available stenght of medication is=125 mg/5ml means 25mg /ml

Q.23. How many kg the weight of patient=41/2.2=18.6 kg

Q.24. per kg required is 5mg

hence for 18.6kg required wwill be=18.6x5=93mg

it is seen that 25mg is in 1 ml

1 mg will be supplied by =1/25 ml

Hence required 93 mg will be supplied by the=(1/25)x93 ml=3.7 ml